OpenStudy (anonymous):

1. Evaluate the function f(x) = 2x + 2, when x = –1. (Points : 1) 4 –4 3 0

OpenStudy (anonymous):

Please help. I have very limited time to get this done...

OpenStudy (nurali):

f(x) = 2x + 2 f(-1)= 2(-1)+2 f(-1)=-2+2 f(-1)=?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Oh my God all you do is plug in the number. Jeez, do you know math around here?????

OpenStudy (anonymous):

@Nurali hiiiiiiiiiiiiiiiiiiiii

OpenStudy (anonymous):

@7123048509388340984308-_-

OpenStudy (nurali):

Wrong answer @bvblover83014

OpenStudy (anonymous):

Can you explain to me what I did wrong??

OpenStudy (nurali):

-2+2=0